# The initial value problem in cosmology: An alternative derivation for beginners

### v1

@@ Abstract
In this article we address the puzzle related to the existence of a proper cosmological solution, where the number of equations exceeds the number of unknown functions. Although only two of the various dynamical equations are independent, the solution of those two equations must have to satisfy all the dynamical equations simultaneously at all instants. We show that this last requirement demands a constraint on the initial condition. When the initial constraint is satisfied, then and only then, one can have proper cosmological dynamics. The whole exercise turns out to be an alternative derivation of the initial value problem in cosmology.
@@ 1 Introduction
A first course in cosmology is rarely preceded by an advanced course in general relativity (GR). In an advanced course in general relativity, one first learns about time like, spacelike or null hypersurfaces. The geometry of hypersurfaces naturally demands the concepts of the first fundamental form and the second fundamental form of the hypersurfaces. Based on the last two conceptions one then proceeds toward the initial value problem in cosmology where the initial conditions are given on a spacelike 3-dimensional hypersurface. To understand the formulation of the initial value problem of the Einstein equation, in the cosmological backdrop, one requires to have a fair amount of understanding of differential geometry and advanced tensor analysis. There are some very good books that can be used by the advanced readers to get hold of these subtle issues [1, 2]. Most of the introductory books on cosmology, as in Ref. [3, 4, 5] do not address these issues as these issues are not that important for understanding cosmological models and the various phenomenological aspects of cosmology. The other books which primarily focus on gravitation, as in Ref. [6, 7], and discuss cosmology as purely an application of GR does not address these issues in any detail. There are a good number of papers related to the dynamics of the Friedmann-Lemaitre-Robertson-Walker (FLRW) model in cosmology, as Refs. [8, 9, 10, 11, 12], where the authors discuss various aspects of cosmology. These sets of papers are written for beginners and are both informative as well as quantitative. None of these literature does discuss the initial value problem of the Einstein equations. Most of the authors comment on the interdependence of the various dynamical equations and then start to analyze the various cosmological models resulting from various kinds of matter present in the universe. This is a perfect way to start and the authors cannot be blamed for not including the initial value problem because the initial value problem, for the case of Einstein equation, is more mathematics-oriented. A good discussion on this issue can be obtained in Ref. [13], but this reference is not suitable for readers who have not taken a course in differential geometry and topology. As a result of the complexity arising from the very nature of the Einstein equation, the initial value problem remains untouched by beginners who start to understand cosmology from university courses. In this article, we will try to address this initial value problem, in the cosmological case. Although the problem is mathematical in nature we will try to present the substance in such a way so that even students who have some knowledge of the Friedmann equations will be able to grasp the main points.
In FLRW cosmology it is known that there are multiple dynamical equations that are obtained by various kinds of combinations of the two Einstein equations. The combination of the two equations may not always be purely algebraic, one may need to differentiate one of the equations and add the result suitably with the other. As a result of these combinations, one can obtain four different equations (or more) [^1] The number of equations may depend upon the scheme in which the two Einstein equations are manipulated and various authors work with various kinds of those schemes. In this article we will primarily focus on four equations. The arguments presented in this article can be applied for any other schemes. out of which only two are independent. It turns out that there are two unknown functions: one is the scale-factor appearing in the FLRW metric and the other is the energy density of matter. Here we are assuming that the fluid representing matter has a barotropic equation of state. In such a case we have more number of equations than unknowns. Although the four equations are not independent, still they are different equations. For a proper solution of the Einstein equation, all of the four differential equations must have to agree simultaneously for all relevant time instants when the actual dynamics is produced from only two differential equations. We will show that a unique dynamics is only obtained when one chooses the initial condition judiciously. The initial conditions, in the case of the cosmological solutions, are constrained by one of the Einstein equations. This constraint plays a vital role in cosmological dynamics. The results we obtain are similar to the one described in Ref. [1] and as a result, our work can be interpreted as an alternative derivation of the cosmological initial value problem. Although the initial value problem in GR is primarily a mathematical problem, it has immense importance to the cosmologists who routinely work with Friedmann equations. If one is not fully aware of the initial value problem one can easily produce wrong results unknowingly.
The initial value problem for the Einstein equation has been worked out previously and we do not claim to present any original new result. We claim that our work is a new alternative derivation of the cosmological initial value problem which can be appreciated by anyone who has just started to know cosmology and has taken an introductory course in GR. The standard derivation of the cosmological initial value problem is inherently more geometric in nature where the initial conditions are specified on a spacelike 3-dimensional Cauchy surface. The effect of the (3+1)-dimensional Einstein equation produces constraints on the initial value on the Cauchy surface. The standard derivation is mathematically rigorous and complete. It has only one disadvantage, it cannot be accessed by students who have not taken some courses on differential geometry or advanced GR. In contrast, our alternative derivation is less geometric in nature. In deriving our results we use the continuity of the cosmological solution and time symmetry of the Einstein equation. Throughout this article, we have emphasized more on intuition. We think the new students who have just started to learn Friedmann equations will be able to follow the whole article.
The material in the article is presented in the following way. In the next section, we present the cosmological solution of the Einstein equation. In this section, we have not emphasized the derivation of the equations as we assume the reader to know these results. This section primarily sets the convention followed in the rest of the article. In section 3, we have posed the question related to the dynamics emanating from the Einstein equation. Section 4 tries to answer the previous question for ever-expanding or ever-contracting spacetimes. In this section, it is seen that the results may depend upon the existence of a turning point, where the time derivative of the scale-factor of the FLRW spacetime changes sign. In section 5 we address the problems related to the turning points. The last section concludes the article with a brief discussion of the results.
@@ 2 A single barotropic fluid in FLRW spacetime and the Einstein equation
Let us start from the beginning. We start with the Einstein equation in general relativity (GR):
$G_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi GT_{\mu\nu}+\Lambda g_{\mu\nu}\,,$ (1)
where $g_{\mu\nu}$ is the metric tensor, $R_{\mu\nu}$ is the Ricci tensor, $R=g^{\mu\nu}R_{\mu\nu}$ is the Ricci scalar, $T_{\mu\nu}$ is the energy-momentum tensor for matter, $\Lambda$ is the cosmological constant and $G$ is the universal gravitational constant. In our units, the speed of light $c=1$. The affine connection coefficients are specified as
$\Gamma^{\mu}_{\alpha\beta}=\frac{1}{2}g^{\mu\sigma}(\partial_{\alpha}g_{\sigma\beta}+\partial_{\beta}g_{\sigma\alpha}-\partial_{\sigma}g_{\alpha\beta})\,.$ (2)
The Riemann tensor is given by
$R^{\mu}_{\,\,\,\,\nu\alpha\beta}=\partial_{\alpha}\Gamma^{\mu}_{\nu\beta}-\partial_{\beta}\Gamma^{\mu}_{\nu\alpha}+\Gamma^{\mu}_{\sigma\alpha}\Gamma^{\sigma}_{\nu\beta}-\Gamma^{\mu}_{\sigma\beta}\Gamma^{\sigma}_{\nu\alpha}\,.$ (3)
Here we are using the Einstein summation convention and any repeated index is summed up. In the above equation we have two summations on $\sigma$. Henceforth we will always use this summation convention. We get the Ricci tensor from the Riemann tensor by contraction of suitable indices as
$R_{\nu\beta}\equiv R^{\mu}_{\,\,\,\,\nu\mu\beta}=\partial_{\mu}\Gamma^{\mu}_{\nu\beta}-\partial_{\beta}\Gamma^{\mu}_{\nu\mu}+\Gamma^{\mu}_{\sigma\mu}\Gamma^{\sigma}_{\nu\beta}-\Gamma^{\mu}_{\sigma\beta}\Gamma^{\sigma}_{\nu\mu}\,.$ (4)
Standard textbooks on GR specifies that due to the Bianchi identity, followed by the Riemann tensor, one will always have
$\nabla_{\mu}G^{\mu\nu}=0\,,$ (5)
where the covariant derivative is defined as
$\nabla_{\mu}G^{\mu\nu}=\partial_{\mu}G^{\mu\nu}+\Gamma^{\mu}_{\mu\alpha}G^{\nu\alpha}+\Gamma^{\nu}_{\mu\alpha}G^{\mu\alpha}\,.$ (6)
The above result is important as it sets an important condition on the energy-momentum tensor for matter.
In GR the affine connection coefficients are symmetric in the lower indices. The Ricci tensor and the metric tensor are also symmetric second rank tensors. This implies $G_{\mu\nu}$ is a symmetric tensor and consequently $T_{\mu\nu}$ should also be a symmetric tensor. From Eq. (5) we immediately get
$\nabla_{\mu}T^{\mu\nu}=0\,.$ (7)
The original Einstein equation remains consistent as because in GR the metric tensor is covariantly constant: $\nabla_{\mu}g^{\mu\nu}=0$. It is seen that the conservation of the energy-momentum tensor is a condition arising from the structure of the Einstein tensor.
The spacetime metric is chosen to be the isotropic and homogeneous FLRW spacetime:
@@
where $a(t)$ is the scale-factor for the expanding or contracting spacetime and $k$ is a constant which can take values ±$\pm 1$ or 00. The spatial coordinate $r$, used in the above line element is called a comoving coordinate. One must be careful in interpreting the above line element. Previously we have said that we are using units where speed of light is one. As a consequence length and time have the same dimension. We can interpret $ds^{2}$ having a dimension of squared length. In this unit what will be the dimensions of $r$, $a(t)$ and $k$? When $k=0$, $r$ does have the dimension of length and $a(t)$ can be dimensionless. When $k=\pm 1$ and dimensionless, $r$ cannot carry length dimension. In such cases $a(t)$ carries the dimension of length. As $k$ comes out from the geometry of the FLRW spacetime, we will like to keep it dimensionless.
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose energy momentum tensor is given by
$T_{\mu\nu}=(\rho+P)u_{\mu}u_{\nu}+Pg_{\mu\nu}\,,$ (9)
where $\rho$ is the energy density of the fluid, $P$ is its pressure and the fluid ‘elements move’ with a time like 4-velocity $u_{\mu}=dx^{\mu}/d\tau$ where
$u_{\mu}u^{\mu}=-1\,.$
In FLRW coordinates the proper time, $\tau$, is the same as the coordinate time $t$ itself. In the FLRW metric, in the comoving coordinate, we will have $u^{\mu}=(1,{\bf 0})$ where ${\bf 0}$ specifies that the fluid elements do not move spatially. For a barotropic fluid one must also have to specify the equation of state (EoS):
@@
where $\omega$, which is a dimensionless quantity in our system of units, is often called the EoS. The EoS is a constant. For a perfect fluid the only nonzero components of the energy-momentum tensor are
$T_{00}=\rho\,,\,\,\,\,\,\,T_{ij}=Pg_{ij}\,.$ (11)
Written in this form one knows how the energy-momentum tensor components are related with the energy density and pressure of the fluid.
In FLRW spacetime, the nonzero components of the Einstein tensor are $G_{00}$ and $G_{ij}$ and these tensor components can be equated to the only nonzero components of the energy-momentum tensor $T_{00}$ and $T_{ij}$. The relevant equations are $G_{00}=8\pi GT_{00}+\Lambda g_{00}$ and $G_{ij}=8\pi GT_{ij}+\Lambda g_{ij}$. These two equations yield
@@
Except the above three equations, which govern the dynamics of the system, there is another one. This equation comes from the energy-momentum conservation condition given in Eq. (7). In FLRW spacetime this condition becomes
@@
As because this equation is an outcome of the fact that the left hand side of the Einstein equation satisfies the condition arising from Bianchi identity and is true because of the particular form of the Einstein tensor, this equation can be derived from the first Friedmann equation and the second Friedmann equation. We will call Eq. (15) as the continuity equation in this work. As like the continuity equations in other branches of physics, this equation also says how the energy density of the system continuously adjusts itself so that the net energy-momentum of the system is conserved.
@@
Now we have four dynamical equations for the development of the FLRW spacetime. We emphasize here that, although we know that we have two independent differential equations coming out from the Einstein equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved. The reason being, none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. We have three functions to determine: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS we know $P(t)$ can be expressed in terms of the energy density and hence we have only two unknown functions. How do we figure out the development of this dynamical system? Of course all four of these equations are not independent of each other, only two of these four equations must be enough to determine the dynamical evolution of the system. How to choose which two of them are really relevant? We will answer this question in a subsequent section. In the next section we specify the problem in its full generality.
3 Setting up the dynamical system
To specify the problem in its full generality we first give a simple definition of a dynamical system. A dynamical system consists of a set of differential equations and a specific initial configuration. Given these two things the dynamical system is expected to predict the time development of the system. In our case we have a dynamical system at hand. Our dynamical system consists of four ordinary, nonlinear, coupled differential equations of different order and a set of initial conditions. The problem is that, we have four differential equations and we require only two out of them and consequently two of the equations will be left out from our system. Suppose the dynamical system evolves from time $t_{i}$ to $t_{f}$ via the joint solution of the two chosen differential equations. What will happen to the other two differential equations which were left out? If the system is not an overdetermined system, consistency requires that the solutions $a(t)$ and $\rho(t)$, during evolution from $t_{i}$ to $t_{f}$, must also satisfy the two equations which are left out during the same time interval. As all the four equations available to us are valid differential equations expressing some physical content, but may not be independent of each other, all of them must agree for any arbitrary time, $t$, in the interval $t_{i}\leq t\leq t_{f}$. One has to be a bit careful to show the agreement as all the differential equations are not of the same degree, some are first order differential equations and some are second order differential equations. As the various differential equations are not of the same degree one has to find out a way to specify equivalent initial conditions for all the differential equations. We will discuss the point of equivalent initial conditions later.
@@ 3 Setting up the dynamical system
We can now be more specific about the problem of time evolution in the FLRW system. Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $[t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set of initial conditions, all the four equations will never match and the system will behave as an overdetermined system. In the last case we have to conclude that the Einstein equation does not yield any specific solution. We present various cases to illustrate the various possibilities in the present problem.
@@ 4 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. We will relax the last assumption in the next section and show that for intervals where we have $\dot{a}(t_{0})=0$, at some point $t_{0}$ in the interval, we can still obtain specific solutions of the Einstein equation. To present our case systematically we present a chronological description of the dynamics produced by various pairs of the relevant equations. We will first try to see whether we can reproduce the differential equations left out from our chosen pair of equations. Although the four differential equations are produced essentially from two independent dynamical equations, it will be seen that one may not get back all the equations from any pair of relevant equations. We will then try to see whether the solution produced from the chosen pair of equations simultaneously satisfy the other equations.
Before we start we must point out that the time interval $[t_{i}\,,\,t_{f}]$ has to be chosen carefully. We have demanded the interval to be such that the scale-factor always increases or decreases with time inside the interval. We have tacitly assumed that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
@@ 4.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. We will show that if $\dot{a}\neq 0$, then in general the pressure equation is obtained from our chosen equations. When we solve the Friedmann equation and the continuity equation we do not explicitly require the expression of $\ddot{a}$. If we try to figure out the value of $\ddot{a}$ from our chosen equations, we exactly get the pressure equation.
To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
@@
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $[t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. Now one can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.

### v2

# Redundancy of the cosmological evolution equations and its relationship with the initial conditions

@@ Abstract
It is known that in Friedmann-Lemaitre-Robertson-Walker cosmology one has more number of dynamical equations, compared to the number of unknown variables. This fact makes some equations redundant. The situation becomes complicated because all the relevant differential equations in cosmology are not of the same order. In this article we study the fate of the redundant equations. We show that this redundancy is inevitable in general relativity and a modern modification of gravity, called $f(R)$ gravity. It is shown that this redundancy predicts a special role to one of the Friedmann equations, which constrains the initial values of the problem. It is shown that the initial constraint condition itself evolves with time. Our method of analyzing the dynamical structure of the theories relies on an operational approach and can be generalized further.
@@ 1 Introduction
The basics of modern cosmology are taught at the masters level in various institutions. The standard paradigm of introductory cosmology is well addressed in Refs. [1, 2, 3] whereas Refs. [4, 5, 6] focusses more on general relativity (GR). These books give a general understanding based on which someone can build up his/her conceptions in cosmology. Readers who will like to know more about the various aspects of modern cosmology based on the Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime can consult the excellent pedagogical articles in Refs. [7, 8, 9, 10, 11]. While going through the basics of cosmology one first learns to tackle the dynamical equations in cosmology, commonly called the Friedmann equations. These Friedmann equations are nonlinear ordinary differential equations which originate from the Einstein equations in GR which are second order nonlinear partial differential equations in three spatial and one time dimensions. While dealing with these equations one first observes that even in the simplest situation, where the universe only contains one barotropic fluid, there is a problem of redundancy. There are simply more number of cosmological evolution equations than required. The situation is more challenging as all the available differential equations are not of the same order. Naturally the question arises why such a problem appears at all and how are we going to obtain proper cosmological development from these equations in a consistent manner. In this article we will primarily be addressing this redundancy problem in cosmology. We will first show that this redundancy problem is intimately related to the initial value problem in cosmology. It turns out that any arbitrary initial condition will not produce proper cosmological dynamics, only initial conditions satisfying a particular constraint will produce proper dynamics. Then we show that the redundancy problem in cosmology has its own logic in GR. We will also briefly state how our analysis in GR can be generalized in a more modern variation of a theory of gravity, called $f(R)$ gravity.
The article is intended for readers who have taken a basic course of introductory cosmology and knows how to solve the Friedmann equations in standard scenarios: As in matter domination or in radiation domination [^1] We assume that the intended reader to be aware of at least some of the references cited in the last paragraph.. What has been presented in this article has deeper connections to the theory of second order nonlinear partial differential equations, involving time development, in three dimensional space.
The interesting point which becomes apparent from the present article is related to the fact that we can learn many interesting properties related to cosmological development in GR, as well as in $f(R)$ gravity, from an operational approach where our aim is solely to solve the dynamical equations unambiguously. Such an approach ultimately leads to a thorough understanding of the logical structure of the theories.
@@ 2 A single barotropic fluid in FLRW spacetime and the Einstein equation
The spacetime metric, in cosmology, is the isotropic and homogeneous FLRW spacetime:
@@
where $a(t)$ is the scale-factor for the expanding or contracting spacetime and $k$ is a constant which can take values $\pm 1$ or 0. If $k=0$, the curvature of the spatial hypersurfaces in FLRW spacetime is zero, indicating flat spatial hypersurfaces. On the other hand k=+1,1k=+1,\,-1, specifies that the spatial hypersurfaces are positively curved (as on the surface of a sphere) or negatively curved (as in a saddle), respectively. The spatial coordinate $r$, used in the above line element is called a comoving coordinate. One must be careful in interpreting the above line element. We are using units where speed of light is one. As a consequence length and time have the same dimension. We can interpret $ds^{2}$ having a dimension of squared length. In this unit what will be the dimensions of $r$, $a(t)$ and $k$? When $k=0$, $r$ does have the dimension of length and $a(t)$ can be dimensionless. When $k=\pm 1$ and dimensionless, $r$ cannot carry length dimension. In such cases $a(t)$ carries the dimension of length. As $k$ comes out from the geometry of the FLRW spacetime, we will like to keep it dimensionless.
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose equation of state (EoS):
@@
where $\omega$, which is a dimensionless quantity in our system of units, is often called the EoS parameter. Here $\rho$ is the energy-density of the fluid and $P$ specifies its isotropic pressure. The EoS parameter is a constant. In presence of a barotropic fluid, the Einstein equations of FLRW cosmology are:
@@
Except the above three equations, which govern the dynamics of the system, there is another one. This equation comes from the energy-momentum conservation condition. The energy-momentum conservation condition in GR is an outcome of the Bianchi identity. In FLRW spacetime this condition becomes
@@
We will call Eq. (6) as the continuity equation in this work. As like the continuity equations in other branches of physics, this equation also says how the energy density of the system continuously adjusts itself so that the net energy-momentum of the system is conserved.
@@
Now we have four dynamical equations for the development of the FLRW spacetime. We emphasize here that, although we know that we have two independent differential equations coming out from the Einstein equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved. The reason being, none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. We have three functions to determine: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS we know $P(t)$ can be expressed in terms of the energy density and hence we have only two unknown functions. How do we figure out the development of this dynamical system?
@@
We can now be more specific about the problem of time evolution in the FLRW system. Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set, all the four equations will never match and the system will behave as an overdetermined system. In the last case we have to conclude that the Einstein equation does not yield any specific solution. We present various cases to illustrate the various possibilities in the present problem.
@@ 3 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. We will relax the last assumption in the next section and show that for intervals where we have $\dot{a}(t_{0})=0$, at some point $t_{0}$ in the interval, we can still obtain specific solutions of the Einstein equation. We will first try to see whether we can reproduce the differential equations left out from our chosen pair of equations. Although the four differential equations are produced essentially from two independent dynamical equations, it will be seen that one may not get back all the equations from any pair of relevant equations. We will then try to see whether the solution produced from the chosen pair of equations simultaneously satisfy the other equations.
Before we start we must point out that the time interval $t_{i}\,,\,t_{f}]$ has to be chosen carefully. We have demanded the interval to be such that the scale-factor always increases or decreases with time inside the interval. We have tacitly assumed that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
@@ 3.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
@@
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. Now, one can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.